# Delay-Multiply-And-Sum Beamforming for Real-Time In-Air Acoustic Imaging

### v1

@@ III-B Real-World Performance
A dataset was recorded using an in-air 3D sonar called eRTIS[laurijssen2025ruggedizedultrasoundsensingharsh], mounted on a mobile robot driving around. It features a 32-element MEMS microphone array and a broadband capacitive transducer, emitting a $2.5\text{\,}\mathrm{ms}$ broadband chirp call between $25\text{\,}\mathrm{kHz}$ and $50\text{\,}\mathrm{kHz}$. These measurements were then used to create acoustic images with all beamforming methods, from DAS to DMAS5, and with and without the optional coherence factor post-processing. Beamforming was performed in 3601 directions with their azimuth angle ranging between -$90\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, with an angular spacing of $0.05\text{\,}\mathrm{\SIUnitSymbolDegree}$ to create a horizontal scan image at elevation angle $0\text{\,}\mathrm{\SIUnitSymbolDegree}$. The recorded signals had 163 840 samples ($0.0364\text{\,}\mathrm{s}$ at $450\text{\,}\mathrm{kHz}$). All images were constructed using the same matched filter and envelope detection using a low-pass filter at 5kHz[jansen2019gpu].
A single frame of this dataset is shown in Fig. 7, and a full video of the dataset with all frames processed is available as supplementary material and online[dmascosysyoutube2025]. From these images, the increased dynamic range of the DMAS techniques can be observed when compared to the naive DAS implementation as well as the coherence-factor weighing increasing the contrast of the image even further as expected from the literature.
@@
TABLE I: Beamforming - Average Execution Times for 1000 Directions
@@
The average timing results for 1000 directions can be found in Table I, where we also show the relative increase when increasing the order as a percentage. In Fig. 8a these same results are shown graphically, specifically for the middle-range GPU platform Jetson Orin Nano, visualizing the differences between the beamforming algorithms. Furthermore, Fig. 8b shows the average execution times across all platforms, specifically for DMAS5-CF. These results clearly show the linear correlation between execution time and amount of directions.
[…]
TABLE II: Coherence Factor - Average Execution Times for 1000 Directions
@@ IV Conclusion
The simulated and real-world results conclusively demonstrate that non-linear beamforming is a powerful tool for high-quality in-air acoustic imaging. By incorporating the DMAS principle, and further extending it through higher-order formulations, we successfully overcome the primary limitations of conventional beamforming, which is the limited dynamic range in the image due to the relatively high sidelobe energy. In-air acoustic sensing has a high dynamic range, which requires a high dynamic range in the imaging function to yield acoustic images that are useful in practical applications. In addition, the coherence factor post-processing step additionally improves the image contrast when applied to the beamformer outputs. The key contribution of this non-linear approach is the significant suppression of incoherent signals and sidelobe artifacts, leading to a substantial enhancement in image contrast across real-world, broadband measurements.
@@
In future work, we want to further explore the larger design space of acoustic imaging systems, taking into account different array geometries, and quantify the mutual influence between the array layout, beamformer operation and resulting image parameters. Furthermore, we wish to validate the DMAS-class methods against traditional beamforming techniques such as MVDR and MUSIC, which was out of scope for this paper. Indeed, as stated before, classic powerhouses in the array signal processing literature such as MVDR and MUSIC require multiple snapshots and operate in a narrowband signal model, which are not applicable in in-air sonar sensing, where large signal bandwidths and sparse or irregular array topologies have shown to be very beneficial for practical imaging applications.

### v2

@@ III-B Real-World Performance
A dataset was recorded using an in-air 3D sonar called eRTIS[13], mounted on a mobile robot driving around. It features a 32-element MEMS microphone array and a broadband capacitive transducer, emitting a $2.5\text{\,}\mathrm{ms}$ broadband chirp call between $25\text{\,}\mathrm{kHz}$ and $50\text{\,}\mathrm{kHz}$. These measurements were then used to create acoustic images with all beamforming methods, from DAS to DMAS5, and with and without the optional coherence factor post-processing. Beamforming was performed in 3601 directions with their azimuth angle ranging between -$90\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, with an angular spacing of $0.05\text{\,}\mathrm{\SIUnitSymbolDegree}$ to create a horizontal scan image at elevation angle $0\text{\,}\mathrm{\SIUnitSymbolDegree}$. The recorded signals had 163 840 samples, with a duration of $0.0364\text{\,}\mathrm{s}$ at $450\text{\,}\mathrm{kHz}$. All images were constructed using the same matched filter and envelope detection using a low-pass filter at 5kHz[7]. A single frame of this dataset is shown in Fig. 7, and a full video of the dataset with all frames processed is available as supplementary material and online[6].
From these images, the increased dynamic range of the DMAS techniques can be observed when compared to the naive DAS implementation as well as the coherence-factor weighing increasing the contrast of the image even further as expected from the literature.
@@
TABLE II: Beamforming - Average Execution Times for 1000 Directions
@@
TABLE III: Coherence Factor - Average Execution Times for 1000 Directions
@@ IV Conclusion
While the proposed DMAS-CF method offers superior contrast, it is important to acknowledge the trade-offs inherent to non-linear multiplicative beamforming. These methods are known to be sensitive to off-axis clutter, which can produce artifacts due to non-linear cross-terms. In the context of in-air acoustic imaging, the impact of these artifacts is generally less pronounced because the imaging scenes are typically sparse—comprising a limited number of discrete sources or reflectors rather than the dense, continuous scattering as seen, for example, in the biomedical field when imaging biological tissue. Furthermore, because artifacts resulting from the multiplication of uncorrelated off-axis clutter lack spatial coherence across the array, the proposed coherence factor acts as a spatial filter that effectively suppresses these non-linear artifacts while preserving the main-lobe integrity. The simulated and real-world results conclusively demonstrate that non-linear beamforming is a powerful tool for high-quality in-air acoustic imaging.
By incorporating the DMAS principle, and further extending it through higher-order formulations, we successfully overcome the primary limitations of conventional beamforming, which is the limited dynamic range in the image due to the relatively high sidelobe energy. In-air acoustic sensing has a high dynamic range, which requires a high dynamic range in the imaging function to yield acoustic images that are useful in practical applications. In addition, the coherence factor post-processing step additionally improves the image contrast when applied to the beamformer outputs. The key contribution of this non-linear approach is the significant suppression of incoherent signals and sidelobe artifacts, leading to a substantial enhancement in image contrast across real-world, broadband measurements.
@@
In future work, we aim to further explore the larger design space of acoustic imaging systems, taking into account different array geometries to quantify the mutual influence between layout, beamformer operation, and resulting image parameters. Furthermore, we wish to validate DMAS methods against traditional techniques like MVDR and MUSIC. However, these established methods generally require multiple snapshots and narrowband models, making them less applicable to in-air sonar sensing where large signal bandwidths and sparse topologies are beneficial for practical imaging.